# Slot State Space Models

### v5

@@ Abstract
Recent State Space Models (SSMs) such as S4, S5, and Mamba have shown remarkable computational benefits in long-range temporal dependency modeling. However, in many sequence modeling problems, the underlying process is inherently modular and it is of interest to have inductive biases that mimic this modular structure. In this paper, we introduce SlotSSMs, a novel framework for incorporating independent mechanisms into SSMs to preserve or encourage separation of information. Unlike conventional SSMs that maintain a monolithic state vector, SlotSSMs maintains the state as a collection of multiple vectors called slots. Crucially, the state transitions are performed independently per slot with sparse interactions across slots implemented via the bottleneck of self-attention. In experiments, we evaluate our model in object-centric video understanding, 3D visual reasoning, and video prediction tasks, which involve modeling multiple objects and their long-range temporal dependencies. We find that our proposed design offers substantial performance gains over existing sequence modeling methods. Project page is available at https://slotssms.github.io/
@@ 2 Preliminaries
A state space model (SSM) defines a sequence-to-sequence mapping between the input $\bm{e}_{1:T}\in\mathbb{R}^{T\times D}$ and the output $\bm{y}_{1:T}\in\mathbb{R}^{T\times D}$ by the following recurrence [18, 17, 35, 29]:
@@
Here, $T$ denotes the sequence length, $\bm{e}_{t},\bm{y}_{t}\in\mathbb{R}^{D}$ are the input and output vectors at each time step $t$, and $\bm{h}_{t}\in\mathbb{R}^{H}$ is the hidden state that summarizes relevant information from the history $\bm{e}_{\leq t}$. The matrices $\bm{\overline{A}}_{t}\in\mathbb{R}^{H\times H}$, $\bm{\overline{B}}_{t}\in\mathbb{R}^{H\times D}$, and $\bm{C}_{t}\in\mathbb{R}^{D\times H}$ are constructed from learnable parameters of the SSM in specific ways that encourage modeling long-range dependencies while maintaining computational efficiency. For example, $\bm{\overline{A}}_{t}$ commonly takes a diagonal or block-diagonal form, with its (complex) eigenvalues distributed close to the unit circle at initialization [15, 17, 19, 16, 35, 29].
When the SSM is time-invariant, meaning $\bm{\overline{A}}_{t},\bm{\overline{B}}_{t},\bm{C}_{t}$ are constant with respect to $t$, the computation of $\bm{y}_{1:T}$ can be parallelized given $\bm{e}_{1:T}$, leading to efficient training. Recent works [14, 5] have further shown that it is possible to condition $\bm{\overline{A}}_{t},\bm{\overline{B}}_{t},\bm{C}_{t}$ on the input $\bm{e}_{t}$ without losing training efficiency. As illustrated in Figure 2 (Left), they employ learnable functions $\bm{\overline{A}}\colon\mathbb{R}^{D}\to\mathbb{R}^{H\times H},\bm{\overline{B}}\colon\mathbb{R}^{D}\to\mathbb{R}^{H\times D},\bm{C}\colon\mathbb{R}^{D}\to\mathbb{R}^{D\times H}$ to generate input-dependent matrices $\bm{\overline{A}}_{t},\bm{\overline{B}}_{t},\bm{C}_{t}$:
@@
This brings the ability to selectively emphasize or ignore certain information based on the input, leading to more flexible sequence modeling.
Because the (block-)diagonal structure of $\bm{\overline{A}}_{t}$ limits information propagation across dimensions, SSMs are typically interleaved with mixing layers such as linear projections or MLPs to mix information across dimensions. Mixing can also be promoted by using dense $\bm{\overline{B}}_{t}$ and $\bm{C}_{t}$ matrices.
@@ 4 Modular Sequence Modeling with SlotSSM
Figure 3: Sequence modeling architecture with SlotSSM. SlotSSMs are hybrid models combining SSMs and attention mechanisms. Each layer consists of a Slot Encoder, SlotSSM, and Slot Mixer. The Slot Encoder uses a Transformer with cross-attention to extract slots from inputs. The SlotSSM independently updates these slots with separate state transitions. The Slot Mixer introduces interactions between slots through self-attention.
@@ 4.1 Slot Encoder
For example, image inputs can be represented as CNN feature maps ($M$ is the number of cells in the feature map), or as embeddings of non-overlapping image patches ($M$ is the number of patches), as proposed in ViT [7]. To extract $K$ slot representations from $\bm{x}_{t}$, we use $K$ learnable CLS tokens $\{\text{CLS}_{t}^{k}\in\mathbb{R}^{D_{x}}\}_{k=1}^{K}$ as queries and perform cross-attention with the input tokens through a Transformer [40]:
@@ 4.3 Sequence Modeling Architecture
As shown in Figure 3, our proposed architecture stacks the (optional) slot encoder, SlotSSM, and slot mixer together at each layer. We use the subscript ‘$|l$’ to denote the relevant variables at layer $l$. For notational convenience, we denote the output of slot mixer at layer $l-1$ by $\{\bm{x}_{t\mid l-1}^{k}\}_{k=1}^{K_{l-1}}$. This serves as the input to layer $l$. We use $\{\bm{x}_{t\mid 0}^{k}\}_{k=1}^{K_{0}}$ to denote the input in Equation 6, where $K_{0}\coloneqq M$. The computation at each layer $l=1,\dots,L$ can then be summarized as follows:
@@ 6 Related Work
Object-Centric Learning. Object-centric learning seeks to discover modular structures and independent mechanisms [12] such as objects and their relations from multi-object images and videos with weak or no supervision. Recent works are predominantly based on the Slot Attention [25] model, which uses a GRU [3] and competitive attention mechanisms to iteratively refine slot representations [32, 34, 23, 8, 45, 33, 22, 46]. However, GRUs and RNNs in general are prone to vanishing gradient issues [30], and the training must be done in a sequential way. These weaknesses render them incapable of scaling up to long-range videos. Our SlotSSMs framework can be specialized to solve object-centric learning tasks. With SSMs at its core, SlotSSMs enjoy parallelizable training and remarkable long-term memory abilities. As a general framework, SlotSSMs can also tackle other tasks such as long-range visual reasoning.
@@ 7.1 Multi-Object Video Prediction
We begin with a multi-object video modeling task to demonstrate the benefit of incorperating modularity into state space.
Dataset and Task. We utilize the bouncing balls video dataset introduced by [39]. This dataset comprises videos of white balls bouncing off each other within an empty background window. Each ball is initialized with random positions, velocities, and masses, which govern the physical behavior of their interactions. The task is to perform conditional video generation, specifically $p({\bf x}_{T+1:T+W}|{\bf x}_{1:T})$. This task is inherently modular as it requires models to maintain memory of each object’s physical attributes and to understand the interaction rules of moving objects.
Experimental Setup. We train the model using teacher-forcing on video sequences of 20 frames with binary cross-entropy loss as the training objective. During testing, we provide $T=10$ context frames and the model performs an autoregressive rollout of $W=20$ frames, using its own predictions. We evaluate model performance using Mean Squared Error (MSE) loss between the predicted and ground truth images.
Models. We employ the SlotSSM architecture described in Section 4.3, and we adopt the Mamba Block [14] as the SSM implementation for both SlotSSM and baseline models. We use the same number of slots for all layers, and thus only apply the Slot Encoder at the first layer. We compare our model against the following baselines:
  1. 1.
    Single State SSM: Shares the same architecture as our model but uses a monolithic state.
  2. 2.
    Single State SSM (Split): A Single State SSM paired with multi-slot encoder and decoder. Slots are concatenated and provided to the SSM, then split into multiple slots for the decoder.
  3. 3.
    RIM [12]: A slot-based RNN model that uses separate RNN weights for each slot and introduces sparse slot update and interaction based on input attention values.
  4. 4.
    Transformer: Vanilla Transformer model with single input embedding at each time step.
  5. 5.
    SlotTransformer: Transformer model with multiple input slots at each time step, analogous to SlotSSMs.
All models share the same encoder and decoder architectures. For the encoder, we use the Transformer encoder described in Section 4.1, with a single CLS token for single-state models. The decoder consists of 3 Transformer layers with self-attention and cross-attention blocks, where patch-wise positional embeddings self-attend and then serve as queries to attend to the slots.
For all slot-based models, we use 6 slots. For RIM, we set $k=4$ for top-k active modules as described in the original paper. We carefully design hyperparameters to match model sizes across baselines, with one exception: RIM inherently requires a larger model size due to separate RNN weights per slot. Additional implementation detials can be found in Appendix B.
@@
Results. Figure 4 presents the comparative performance of our models. SlotSSM outperforms all baselines, including a marginal improvement over SlotTransformer. The substantial performance gap between SlotSSM and Single State SSM highlights the importance of maintaining modular slot states for effective multi-object dynamics learning. The comparison between Transformer and SlotTransformer also corroborates this finding.
Notably, SlotSSM also significantly outperforms Single State SSM (Split) which uses the same modular encoder and decoder. This result highlights that modularity in temporal modeling—the core contribution of SlotSSM—is the most critical factor for the improved performance. Additionally, the RNN-based multi-slot model, RIM, performs better than other single-state baselines but still lags significantly behind SlotSSM and SlotTransformer.
@@ 7.2 Long-Context Reasoning
Blinking Color Balls Benchmark. The benchmark comprises two variants: the Earliest Color and the Most Frequent Color versions. Both variants consist of episodes of image sequence ${\bf x}_{1:T}$, partitioned into context images ${\bf x}_{1:T-1}$ and a target image ${\bf x}_{T}$. For each timestep in ${\bf x}_{1:T-1}$, all bouncing balls are initially colored white. Then, one ball is randomly selected and assigned a non-white color, which is randomly sampled from a set of 5 predefined colors The coloring of balls in the target image ${\bf x}_{T}$ depends on the variant:
  • Earliest Color: Each ball’s color is determined by the earliest non-white color assigned to it in the context sequence. If a ball is never assigned a non-white color, it remains white.
  • Most Frequent Color: Each ball’s color is determined by the most frequently assigned non-white color across all context frames. In case of ties, the earliest color is used. If a ball is never assigned a non-white color, it remains white.
To transform the dataset into a long-range reasoning task, we patchify each context image into non-overlapping patches and provide the flattened patch sequence as input to the model. We present this process in Figure 5(a). Specifically, each context image is divided into $P\times P$ patches, which are then flattened into a sequence of length $P^{2}$. For a context length of $T-1$, this results in a total input sequence length of $L=(T-1)\times P^{2}$. Consequently, the models must identify and track objects from partial views provided by the patches, while simultaneously remembering and counting the color assignments for each object, making the task highly challenging and non-trivial. The final task is to predict the target image given this long sequential input.
Experimental Setup. We evaluate models on Earliest Color with $T=6$ and Most Frequent Color with $T\in\{6,11\}$, and we employ patch sizes of $P\in\{4,8,16\}$. The resulting input sequence lengths are $L\in\{80,160,320,640,1280,2560\}$. Note that Most Frequent Color with $T=11$ setting poses a greater challenge because the increased context length leads to more color assignments, requiring the models to exhibit stronger memorization and reasoning capabilities.
Models. We employ the same encoder, decoder, and the SlotSSM architectures as described in Section 7.1. For slot encoding, each image patch is treated as a small image and processed by the transformer slot encoder. The resulting slots from the last time step are then provided to the transformer decoder to predict the full target image. We compare our SlotSSM against several baselines: Single State SSM, SlotTransformer, and RIM. Additionally, we introduce a novel multi-slot design called SlotRNNs, which shares model weights across the RNNs of each slot and uses self-attention layers between time steps as the slot mixer. SlotRNNs can be viewed as a special case of RIMs, with shared RNN weights across states and dense state updates instead of sparse state updates. Empirically, we find that SlotRNNs exhibit more stable training and improved performance compared to RIMs. For a fair comparison, all slot-based models use 6 slots. We also carefully match the model size as in Section 7.1
@@
Results. Figure 6 presents the target image prediction error. SlotSSM demonstrates clear advantages over Single State SSM, SlotRNN, and RIM across all sequence lengths. For shorter sequences (length 80 and 160), Single State SSM and SlotRNN maintain relatively low error rates. However, their performance degrades significantly for sequences longer than 320 frames. Surprisingly, despite its modular architecture, RIM fails to generalize on any sequence length. We hypothesize that the separate weights for each slot in RIMs hinder optimization, and this issue is partially addressed by our proposed baseline SlotRNNs, which share weights across slots while maintaining modularity.
SlotTransformer performs competitively for sequences up to 640 frames, which benefits from their direct access to all historical inputs. Nonetheless, SlotSSM exhibits superior long-range reasoning capabilities, particularly for sequences of 1280 and 2560 frames, where other models can not run due to memory and computational constraints. Figure 5(b) highlights the computational efficiency of SlotSSM. While SlotTransformer achieves similar inference latency for shorter sequences, their performance quickly declines as the sequence length increases due to its quadratic computational complexity. In contrast, SlotSSM maintains a stable and efficient inference process across all sequence lengths. Note that due to the high memory consumption of SlotTransformer, a batch size of 6 is used to evaluate the inference latency of all models.
We include qulitative comparisons in Appendix A.3 to offer further insights into these models’ strengths and weaknesses.
@@ 7.3 Unsupervised Object-Centric Learning
Datasets. We evaluate the OC-SlotSSM model using the MOVi video dataset [13], specifically focusing on the MOVi-A and MOVi-B subsets. Both datasets contain videos of up to 10 objects moving in a 3D environment. MOVi-B introduces additional complexity compared to MOVi-A by incorporating a wider variety of object types and multi-colored backgrounds.
Tasks. Following the evaluation protocol in prior object-centric learning works [25, 22], we evaluate the models on two downstream tasks: unsupervised object segmentation and attribute prediction. For the segmentation task, we report the FG-ARI and mIoU metrics on video sequences. In the attribute prediction task, we measure the quality of representations by inferring object properties from them. For discrete attributes such as object shape, we report the prediction accuracy, while for continuous attributes like object position, we use the $R^{2}$ metric.
Models. We compare OC-SlotSSM against SAVi [23], an RNN-based object-centric learning approach. Both models use a CNN encoder to extract image features as model input tokens $\{\bm{x}_{t\mid 0}^{m}\}_{m=1}^{M}$, which are processed by their respective attention mechanisms (inverted attention in OC-SlotSSM and slot attention in SAVi) to produce slots. The slots are then used to reconstruct the image and generate segmentation masks for each object using a spatial broadcast decoder, with reconstruction as the training objective. For unsupervised object segmentation, we directly use the object masks obtained during unsupervised training. To evaluate attribute prediction, we match the slots to object IDs using Hungarian matching based on the segmentation masks, and then employ linear heads and 2-layer MLPs to predict discrete and continuous attributes, respectively, where the slots remain frozen.
Results. Figure 7 presents the results. The results demonstrate that OC-SlotSSM consistently outperforms SAVi in unsupervised object segmentation on both MOVi-A and MOVi-B. The qualitative comparison (Figure 7, left) shows that OC-SlotSSM generates masks with tighter object boundaries and fewer object splitting, which also leads to improved attribute prediction accuracy (Figure 7, right).
Furthermore, our experiments reveal that OC-SlotSSM exhibits superior stability during training compared to SAVi, which tends to collapse into a single slot representing the entire scene when trained for an extended period. This collapse is not reflected in the validation loss, so we apply early stopping based on manual inspection. In contrast, OC-SlotSSM does not suffer from this instability, demonstrating its robustness in learning object-centric representations.
@@
CATER Benchmark. CATER consists of 300-frame video episodes of objects moving around in a 3D environment. The movement of the objects can lead to partial occlusions and even complete coverage of smaller objects by larger ones. The primary task is snitch localization, which involves predicting the location of a golden snitch at the final frame. Note that the snitch is always present but may be invisible due to occlusion or coverage by other objects. To solve this task, models must be able to reason about the snitch’s location based on its last known position and the movement of other objects in a 3D space. Success in this task would demonstrate the models’ capacity for complex visual reasoning and their potential for application in real-world dynamic 3D environments.
Experimental Setup. We consider two experiment settings: direct training and pre-training $+$ fine-tuning. In the direct training setting, models are trained end-to-end on the snitch localization task without any auxiliary objectives. In the pre-training $+$ fine-tuning setting, models are first pre-trained on video-only inputs using a reconstruction objective, and then fine-tuned on the task-specific signal.
During pre-training, we randomly sample 32 frames, which are not necessarily consecutive, from the original 300-frame videos as input to the model. For direct training and fine-tuning, we first split the input sequence into 50 non-overlapping segments, each containing 6 frames. Then, from each segment, we randomly select one frame, resulting in a subsampled sequence of 50 frames that spans the entire video duration. At test time, we evenly sample 50 frames from the video by skipping every 6 frames. The final snitch location is quantized into a 6x6 grid, and the problem is formulated as a classification task.
Models. In our investigation, we evaluate the performance of SlotSSM, OC-SlotSSM, Single State SlotSSM, and SlotTransformer. We exclude RNN-based baselines, such as SAVi, as our preliminary experiments reveal that they are unstable when handling long video inputs and prone to collapse to a constant output. For the visual pre-training setting, we employ a spatial broadcast decoder, common to all models, to reconstruct the input images. During downstream training/fine-tuning, we feed the slots from the final time step into a transformer predictor with single CLS token, followed by a linear layer on the output CLS token to predict the snitch’s position.
@@
Results. Table 1 presents the Top-1 and Top-5 prediction accuracy on the CATER Snitch Localization task. Firstly, consistent with our previous findings, SlotSSM outperforms Single State SSM, which demonstrates the importance of modular structure in latent states for reasoning tasks involving multiple objects. When comparing SlotSSM with SlotTransformer, we see a notable difference between the direct training and pre-training settings. In direct training, SlotTransformer surpasses SlotSSM, possibly due to their optimization advantage, as the model can directly access to all previous states which facilitates learning of the task. However, SlotSSM benefits more from the pre-training phase, potentially attributed to the explicit memory capacity enabled by SSM states. Consequently, pre-trained SlotSSMs outperform their SlotTransformer counterparts.
Remarkably, OC-SlotSSM achieves the highest accuracy, outperforming all baselines by a large margin in both direct training and pre-training settings. This performance gain might be attributed to the explicit decomposition of the scene into object-centric representations, which facilitates the reasoning about object properties, relationships, and interactions.
8 Conclusion & Limitations

### v6

@@ Abstract
Recent State Space Models (SSMs) such as S4, S5, and Mamba have shown remarkable computational benefits in long-range temporal dependency modeling. However, in many sequence modeling problems, the underlying process is inherently modular and it is of interest to have inductive biases that mimic this modular structure. In this paper, we introduce SlotSSMs, a novel framework for incorporating independent mechanisms into SSMs to preserve or encourage separation of information. Unlike conventional SSMs that maintain a monolithic state vector, SlotSSMs maintains the state as a collection of multiple vectors called slots. Crucially, the state transitions are performed independently per slot with sparse interactions across slots implemented via the bottleneck of self-attention. In experiments, we evaluate our model in object-centric learning, 3D visual reasoning, and long-context video understanding tasks, which involve modeling multiple objects and their long-range temporal dependencies. We find that our proposed design offers substantial performance gains over existing sequence modeling methods. Project page is available at https://slotssms.github.io/
@@ 2 Preliminaries
A state space model (SSM) defines a mapping between an input sequence $\bm{e}_{1:T}\in\mathbb{R}^{T\times D}$ and an output sequence $\bm{y}_{1:T}\in\mathbb{R}^{T\times D}$ via the recurrence [28, 27, 58, 50]:
@@
Here, $T$ is the sequence length; $\bm{e}_{t},\bm{y}_{t}\in\mathbb{R}^{D}$ are input and output vectors at time $t$; and $\bm{h}_{t}\in\mathbb{R}^{H}$ is the hidden state summarizing the history $\bm{e}_{\leq t}$. The matrices $\bm{\overline{A}}_{t}\in\mathbb{R}^{H\times H}$, $\bm{\overline{B}}_{t}\in\mathbb{R}^{H\times D}$, and $\bm{C}_{t}\in\mathbb{R}^{D\times H}$ are designed with learnable parameters in specific ways that encourage modeling long-range dependencies while maintaining computational efficiency. For example, $\bm{\overline{A}}_{t}$ commonly takes a diagonal or block-diagonal form, with its (complex) eigenvalues distributed close to the unit circle at initialization [25, 27, 29, 26, 58, 50].
When $\bm{\overline{A}}_{t},\bm{\overline{B}}_{t},\bm{C}_{t}$ are time-invariant (constant over $t$), the computation of $\bm{y}_{1:T}$ can be parallelized, enabling efficient training. Recent works [24, 9] further show that conditioning these matrices on the input $\bm{e}_{t}$ does not hinder training efficiency. They employ learnable functions $\bm{\overline{A}}\colon\mathbb{R}^{D}\to\mathbb{R}^{H\times H},\bm{\overline{B}}\colon\mathbb{R}^{D}\to\mathbb{R}^{H\times D},\bm{C}\colon\mathbb{R}^{D}\to\mathbb{R}^{D\times H}$ to generate input-dependent matrices:
@@
This allows the model to selectively emphasize or ignore information based on the input, leading to more flexible sequence modeling.
Due to the (block-)diagonal structure of $\bm{\overline{A}}_{t}$ limiting cross-dimensional information flow, SSMs are typically interleaved with mixing layers (e.g., linear projections or MLPs) to mix information across dimensions. Alternatively, using dense $\bm{\overline{B}}_{t}$ and $\bm{C}_{t}$ matrices can also enhance mixing.
@@ 4 Modular Sequence Modeling with SlotSSM
Figure 3: Sequence modeling with SlotSSM. Each layer includes a Slot Encoder, SlotSSM, and Slot Mixer. The Slot Encoder uses a Transformer to extract slots from inputs. The SlotSSM independently updates the slots via separate state transitions. The Slot Mixer introduces inter-slot interactions through self-attention.
@@ 4.1 Slot Encoder
For example, image inputs can be CNN feature maps ($M$ is the number of cells in the feature map), or as embeddings of non-overlapping image patches ($M$ is the number of patches), as proposed in ViT [12]. To extract $K$ slot representations from $\bm{x}_{t}$, we use $K$ learnable CLS  [^1] Following the tradition of ViT, we use “CLS” to distingish learnable tokens from observation tokens. tokens $\{\text{CLS}_{t}^{k}\in\mathbb{R}^{D_{x}}\}_{k=1}^{K}$ as queries and perform cross-attention with the input tokens through a Transformer [63]:
@@ 4.3 Sequence Modeling Architecture
[…]
@@ 6 Related Work
Object-Centric Learning. Object-centric learning seeks to discover modular structures and independent mechanisms [20] such as objects and their relations from multi-object images and videos with weak or no supervision [3, 22, 38, 23, 15, 14, 16, 7, 45, 35, 41, 37, 6, 44, 11, 1, 64, 68, 53, 34]. Recent works are predominantly based on the Slot Attention [46] model, which uses a GRU [5] and competitive attention mechanisms to iteratively refine slot representations [54, 57, 39, 13, 69, 55, 36, 70]. However, GRUs and RNNs in general are prone to vanishing gradient issues [51], and the training must be done in a sequential way. These weaknesses render them incapable of scaling up to long-range videos. Additionally, hardware parallelization for video object-centric learning has been explored in [56], however, it incurs quadratic cost unlike ours. Our SlotSSMs framework can be specialized to address object-centric learning tasks effectively. By integrating SSMs at its core, SlotSSMs benefit from parallelizable training and possess remarkable long-term memory capabilities. Moreover, as a versatile framework, SlotSSMs are well-suited to tackle other tasks such as long-range visual reasoning.
@@ 7.1 Multi-Object Video Prediction
We begin with a multi-object video modeling task to demonstrate the benefit of incorporating modularity into state space.
Dataset and Task. We utilize the bouncing balls video dataset introduced by [62], which consists of videos of white balls bouncing off each other in an empty window. Each ball has random initial positions, velocities, and masses, governing their interactions. The task is conditional video generation, specifically $p({\bf x}_{T+1:T+W}|{\bf x}_{1:T})$. This task is inherently modular as it requires models to remember each object’s attributes and interaction rules.
Experimental Setup. We train models on 20-frame sequences using teacher-forcing and binary cross-entropy loss. At test time, given $T=10$ context frames, the model autoregressively predicts $W=20$ future frames using its own outputs. Performance is evaluated using Mean Squared Error (MSE) between predicted and ground-truth images.
Models. We employ the SlotSSM architecture described in Section 4.3 We use the same number of slots across layers and apply the Slot Encoder only at the first layer. We compare our model against several baselines: Single State SSM, which shares the same architecture but uses a monolithic state; Single State SSM (Split), which uses a Single State SSM with multi-slot encoder and decoder—slots are concatenated in SSM, then split into multiple slots for the decoder; RIM[20], a slot-based RNN model with separate RNN weights per slot that introduces sparse slot updates and interactions based on input attention values; Transformer, a vanilla Transformer model with a single input embedding per time step; and SlotTransformer, a Transformer model with multiple input slots at each time step.
All models share the same encoder and decoder architectures. The encoder is a Transformer described in Section 4.1, using a single CLS token for single-state models. The decoder consists of three Transformer layers with self-attention for image patches and cross-attention to query the slots. We use six slots for all slot-based models. For RIM, we set $k=4$ for top-$k$ active modules as in the original paper. We carefully match hyperparameters across baselines to ensure comparable model sizes, except for RIM, which inherently requires a larger model due to separate RNN weights per slot. Additional implementation details are in Appendix C.
@@
Results. Figure 4 compares model performances, showing that SlotSSM outperforms all baselines, including a slight improvement over SlotTransformer. The significant gap between SlotSSM and Single State SSM underscores the importance of modular slot states for effective multi-object dynamics learning, as also evidenced by the comparison between Transformer and SlotTransformer. SlotSSM also significantly outperforms Single State SSM (Split), which uses the same modular encoder and decoder, highlighting that modularity in temporal modeling—the core contribution of SlotSSM—is critical for improved performance. While the RIM model performs better than other single-state baselines, it still lags behind SlotSSM and SlotTransformer.
@@ 7.2 Long-Context Reasoning
Blinking Color Balls Benchmark. This benchmark has two variants—Earliest Color and Most Frequent Color—each consisting of image sequences with context images ${\bf x}_{1:T-1}$ and a target image ${\bf x}_{T}$. In each context image, one ball is randomly selected and assigned a non-white color from five options while others remain white. The coloring of balls in the target image ${\bf x}_{T}$ depends on the variant: in Earliest Color, each ball’s color is its earliest assigned non-white color in the context (remaining white if none); in Most Frequent Color, each ball’s color is the non-white color assigned most frequently during the context (ties broken by earliest assignment; remaining white if none).
To create a long-range reasoning task, we further patchify each context image into non-overlapping $P\times P$ patches and flatten them into a sequence of length $P^{2}$, as shown in Figure 5(a). With context length $T-1$, the total input sequence length is $L=(T-1)\times P^{2}$. Models must identify and track objects from partial patch views while remembering and counting color assignments, making the task highly challenging. The final task is to predict the target image given this long sequential input.
Experimental Setup. We evaluate models on Earliest Color with $T=6$ and Most Frequent Color with $T\in\{6,11\}$, using patch sizes $P\in\{4,8,16\}$. This yields input sequence lengths $L\in\{80,160,320,640,1280,2560\}$. The Most Frequent Color with $T=11$ requires stronger memorization and reasoning capabilities due to longer context and more color assignments, .
Models. We employ the same encoder, decoder, and the SlotSSM architectures as in Section 7.1. For slot encoding, each image patch is treated as an image and processed by the slot encoder. The slots from the last time step are provided to the decoder to predict the full target image. We compare our SlotSSM against several baselines: Single State SSM, SlotTransformer, and RIM. Additionally, we introduce a novel slot-based design called SlotRNNs, which shares RNN weights across slots and uses self-attention layers between time steps as the slot mixer. SlotRNNs can be viewed as a special case of RIMs with shared weights and dense state updates. Empirically, SlotRNNs exhibit more stable training and improved performance compared to RIMs. For fair comparison, all slot-based models use six slots, and we carefully match model sizes as in Section 7.1.
@@
Results. Figure 6 shows that SlotSSM outperforms Single State SSM, SlotRNN, and RIM across all sequence lengths. For shorter sequences (80 and 160), Single State SSM and SlotRNN have relatively low error rates but degrade significantly beyond 320 frames. Surprisingly, RIM fails to generalize at any sequence length, likely due to optimization issues from separate weights per slot; our SlotRNN addresses this by sharing weights across slots while maintaining modularity. SlotTransformer performs competitively up to 640 frames. However, SlotSSM demonstrates superior long-range reasoning, especially at 1280 and 2560 frames, where other models cannot run due to memory or optimization constraints. Figure 5(b) highlights SlotSSM’s computational efficiency. SlotTransformer’s inference latency increases rapidly with sequence length due to quadratic complexity, SlotSSM maintains stable and efficient inference across all lengths. Due to SlotTransformer’s high memory usage, we used a batch size of 6 for latency evaluation. Qualitative comparisons in Appendix B.3 provide further insights into the models’ strengths and weaknesses.
@@ 7.3 Unsupervised Object-Centric Learning
Datasets. We evaluate OC-SlotSSM on the MOVi-A and MOVi-B subsets of the MOVi video dataset [21], which contain videos of up to 10 objects moving in a 3D environment. MOVi-B adds complexity over MOVi-A by including a wider variety of object types and multi-colored backgrounds.
Tasks. Following prior object-centric learning works [46, 36], we evaluate models on two downstream tasks: unsupervised object segmentation and attribute prediction. For segmentation, we report FG-ARI and mIoU metrics. For attribute prediction, we measure the quality of representations by inferring object properties: we report prediction accuracy for discrete attributes (e.g., object shape) and $R^{2}$ for continuous attributes (e.g., object position).
Models. We compare OC-SlotSSM to SAVi [39], an RNN-based object-centric learning approach. Both models use a CNN encoder to extract image features as input tokens ${\bm{x}{t\mid 0}^{m}}{m=1}^{M}$, which are processed by their respective attention mechanisms—inverted attention in OC-SlotSSM and slot attention in SAVi—to produce slots. These slots are then used to reconstruct the image and generate per-object segmentation masks via a spatial broadcast decoder, with reconstruction as the training objective. For unsupervised object segmentation, we directly use the object masks obtained during training. For attribute prediction, we match slots to object IDs using Hungarian matching based on segmentation masks, then use linear heads and 2-layer MLPs to predict discrete and continuous attributes, respectively, keeping the slots frozen.
@@
Results. Results in Figure 7 demonstrate that OC-SlotSSM consistently outperforms SAVi in unsupervised object segmentation on both MOVi-A and MOVi-B. The qualitative comparison (Figure 7, left) shows that OC-SlotSSM generates masks with tighter object boundaries and fewer object splitting, which also leads to improved attribute prediction accuracy (Figure 7, right). Furthermore, we empirically found that OC-SlotSSM exhibits superior stability during training compared to SAVi, which tends to collapse into a single slot representing the entire scene when trained long enough. This collapse is not reflected in the validation loss, so we apply early stopping based on manual inspection. In contrast, OC-SlotSSM does not suffer from this instability, demonstrating its robustness in learning object-centric representations.
@@
CATER Benchmark. CATER consists of 300-frame video episodes of objects moving in a 3D environment. The movement can lead to partial occlusions and even complete coverage of smaller objects by larger ones. The primary task is snitch localization—predicting the golden snitch’s location in the final frame. The snitch is always present but may be occluded. Models must reason about its location based on the last visible position and other objects’ movements. Success in this task demonstrates models’ capacity for complex visual reasoning in dynamic 3D environments.
Experimental Setup. We consider two experiment settings: direct training and pre-training $+$ fine-tuning. In direct training, models are trained end-to-end on the snitch localization task. In pre-training $+$ fine-tuning, models are first pre-trained on video inputs using a reconstruction objective, then fine-tuned on the task-specific signal. During pre-training, we randomly sample 32 frames from the 300-frame videos. For direct training and fine-tuning, we split the sequence into 50 non-overlapping segments of 6 frames each, randomly selecting one frame from each to create a 50-frame sequence spanning the entire video. At test time, we evenly sample 50 frames by skipping every 6 frames. The snitch’s final location is quantized into a 6×6 grid, framing the problem as a classification task.
Models. We evaluate the performance of SlotSSM, OC-SlotSSM, Single State SSM, and SlotTransformer. We exclude RNN-based baselines, as our preliminary experiments reveal that they are unstable when handling long video inputs and prone to collapse to a constant output. For the visual pre-training setting, we employ a spatial broadcast decoder to reconstruct the input images. During downstream training/fine-tuning, we feed the slots from the final step to a transformer predictor with single CLS token, followed by a linear layer on the output CLS token to predict the snitch’s position.
@@
Results. Table 1 presents the Top-1 and Top-5 accuracy on the CATER Snitch Localization task. Consistent with our previous findings, SlotSSM outperforms Single State SSM, highlighting the importance of modular latent structures. Comparing SlotSSM with SlotTransformer, we see notable differences between direct training and pre-training settings: in direct training, SlotTransformer surpasses SlotSSM, possibly due to optimization advantages from direct access to all previous states; however, SlotSSM benefits more from pre-training, likely due to the explicit memory capacity of SSM states, consequently, pre-trained SlotSSMs outperforming their SlotTransformer counterparts.
Remarkably, OC-SlotSSM achieves the highest accuracy, outperforming all baselines by a large margin in both direct training and pre-training settings. This performance gain may be attributed to the explicit decomposition into object-centric representations, which facilitates reasoning about object properties, relationships, and interactions.
8 Conclusion